OpenStudy (anonymous):

can someone help me find the solution to: arctan(2sin(pi/3)) a walkthrough would help!

OpenStudy (anonymous):

Use a 30-60-90 right triangle. \[\tan ^{-1}\left(2 \sin \left(\frac{\pi }{3}\right)\right)=\tan ^{-1}\left(\frac{2 \sqrt{3}}{2}\right)=\tan ^{-1}\left(\sqrt{3}\right)=\frac{\pi }{3} \]